mathslover (mathslover):

Show how the Kepler's third planetary motion led Newton to establish the inverse-square rule for gravitation.

mathslover (mathslover):

@hartnn @Algebraic! @ajprincess @experimentX @experimentX

mathslover (mathslover):

oops tagged experiment x two times sorry :D

mathslover (mathslover):

@eyust707

mathslover (mathslover):

well I think I should do something for u all .. should I draw a diagram ?

OpenStudy (anonymous):

LOL this is easy.

OpenStudy (anonymous):

Consider the following image.

mathslover (mathslover):

ok

OpenStudy (anonymous):

https://dl.dropbox.com/u/63664351/Inverse%20Square%20Law.PNG The lines represent the flux emanating from the source. The total number of flux lines depends on the strength of the source and is constant with increasing distance. A greater density of flux lines (lines per unit area) means a stronger field. Since the surface area of a sphere (which is 4πr2 ) is proportional to the square of the radius, as the emitted radiation gets farther from the source, it is spread out over an area that is increasing in proportion to the square of the distance from the source. Hence, the intensity of radiation passing through any unit area (directly facing the point source) is inversely proportional to the square of the distance from the point source.

mathslover (mathslover):

not exactly what I wanted @twitter do you know the eqn for that?

mathslover (mathslover):

\[\large{\frac{r^3}{T^2} \textbf{is constant}}\]

mathslover (mathslover):

indirectly you have to prove that...

OpenStudy (anonymous):

Intensity = C/r^2 ; where C is constant and Intensity is the gravitation? Oops. Looks like I give you the most fundamental one. To get that equation, you just need to manipulate the algebra.

mathslover (mathslover):

or simply let the question be : prove that : \[\large{\frac{r^3}{T^2}\quad \text{is constant}}\]

OpenStudy (anonymous):

From Kepler equation? Can you give me that eqn. I forget.

mathslover (mathslover):

\[\large{F=\frac{mv^2}{R}}\] or \[\large{F \propto \frac{v^2}{R}}\]

OpenStudy (anonymous):

But seriously the imagine I give you before is really fundamental. You can even derive electricity, light, electromagnetic radiation and even gravitation F = Gm1m2/R^2 simply from there!

mathslover (mathslover):

since \(\large {v = \frac{2\pi R}{T}}\)

sam (.sam.):

\[F=\frac{GMm}{r^2} \\ \\ \text{Due \to planets rotating around the sun, }F=\frac{mv^2}{r} \\ \\ \frac{mv^2}{r}=\frac{GMm}{r^2} \\ \\ \frac{\cancel{m}v^2}{\cancel{r}}=\frac{GM\cancel{m}}{r^\cancel{2}} \\ \\ \text{Rotational period, } v=\frac{2 \pi r}{T} \] \[\frac{4\pi^2 r^2}{T^2}=\frac{GM}{r} \\ \\ r^3=\frac{GM}{4 \pi^2}T^2 \\ \\ \frac{GM}{4\pi^2}is~constant, \\ \\ \text{Therefore, } kT^2=r^3\]

OpenStudy (experimentx):

yeah .. this concept is just conversation of gravitational flux i seem to have forgotten how to arrive at this from kepler's third law.

mathslover (mathslover):

hence : \[\large{v \propto \frac{R}{T}}\] \[\large{v^2\propto \frac{R^3}{T^2} \times \frac{1}{R}}\]

mathslover (mathslover):

wow 2 proves ... we can also prove that : \[\large{F \propto \frac{1}{R^2}}\]

mathslover (mathslover):

Thanks @.Sam. :)

OpenStudy (anonymous):

Oops someone done it first

OpenStudy (anonymous):

@experimentX you don't arrive to this from Kepler Law. You derive this to Kepler Law, and other LAWS that have r^2 in it.

OpenStudy (experimentx):

i remember it being very simple ... within in 2-3 steps.

OpenStudy (anonymous):

Really? I get it using pure logic and calculus. From the function of circle to get the Area of a sphere. Voila.

OpenStudy (experimentx):

well probably the same way as .sam did ... just manipulation ... ellipses make things uneasy

OpenStudy (anonymous):

um. my i know your current education level? I meant you derive this to get to other fundamental laws in physics. But for high school you don't need to know this.

OpenStudy (experimentx):

i'm out of university.

OpenStudy (experimentx):

i always wondered how kepler got things without calculus ahead of newton.